# Linear and Sublinear Diversities

### v2

@@ Abstract
Diversities are an extension of the concept of a metric space, where a non-negative value is assigned to every finite set of points, rather than just pairs. A general theory of diversities has been developed which exhibits many deep analogies to metric space theory but also veers off in new directions. Just as many of the most important aspects of metric space theory involve metrics defined on $\mathbb{R}^{k}$, many applications of diversity theory require a specialized theory for diversities defined on $\mathbb{R}^{k}$, as we develop here. We focus on two fundamental classes of diversities defined on $\mathbb{R}^{k}$: those that are Minkowski linear and those that are Minkowski sublinear. Many well-known functions in convex analysis belong to these classes, including diameter, circumradius and mean width. We derive surprising characterizations of these classes, and establish elegant connections between them. Motivated by classical results in metric geometry, and connections with combinatorial optimization, we then examine embeddability of finite diversities into $\mathbb{R}^{k}$. We prove that a finite diversity can be embedded into a linear diversity exactly when it has negative type and that it can be embedded into a sublinear diversity exactly when it corresponds to a generalized circumradius.
@@ 1 Introduction
A diversity [6] is a pair $(X,\delta)$ where $X$ is a set and $\delta$ is a non-negative function defined on finite subsets of $X$ satisfying
@@
As such, diversities are set-based analogues of metric spaces, and in fact the restriction of a diversity to pairs is a metric space [6].
Properties (D1) and (D2) are equivalent to (D1) together with monotonicity
@@
Many well-known set functions are diversities: the diameter of a set; the length of a connecting Steiner tree; the circumradius; the length of a minimal traveling salesperson tour; the mean width; the size of a smallest enclosing zonotope. Two set functions which fail to be diversities are genetic diversity
$\pi(A)=\mbox{$\binom{|A|}{2}^{-1}$}\sum_{a,b\in A}d(a,b),$
which is not monotonic (D3), and volume of convex hull, which fails (D2) and (D4).
There are broad classes of diversities just like there are broad classes of metrics. The $\ell_{1}$ metrics have the form
$d_{1}(a,b)=\sum_{i}|a_{i}-b_{i}|,$
while $\ell_{1}$ diversities [7] have the form
$\delta_{1}(A)=\sum_{i}\max_{a,b\in A}|a_{i}-b_{i}|.$
Negative-type metrics satisfy
$\sum_{a,b}x_{a}x_{b}\,d(a,b)\leq 0$
for all zero-sum vectors $x$ while negative type diversities [27] satisfy
$\sum_{A,B}x_{A}x_{B}\,\delta(A\cup B)\leq 0$
for all zero-sum vectors $x$ with $x_{\emptyset}=0$.
The theory of diversities sometimes runs in parallel with that of metric spaces and other times veers off in new directions. In the first paper on diversities [6] we explored how concepts of hyperconvexity, injectivity and the tight span extended, and to an extent enriched, the analogous metric concepts. In [7] we showed that the ‘geometry of graphs’ [18] linking metric embeddings to approximation algorithms on graphs has a parallel ‘geometry of hypergraphs’ linking diversity embeddings to approximation algorithms on hypergraphs. Jozefiak and Shephard [17] use this approach to obtain the best known approximation algorithms for several hypergraph optimization problems.
Diversities turn out to be an exemplary class of metric structures, exhibiting fascinating connections with model theory and Urysohn’s universal space [4, 5, 15, 14]. Other directions that have been pursued are a diversity analogue of ultrametric and normed spaces [13, 20, 12] and new diversity-based results in fixed point theory [9, 22].
Our focus here is on the intersection of diversity theory, geometry and convex analysis. Recall that the Minkowski sum of two subsets $A,B\subseteq\mathbb{R}^{k}$ is given by $A+B=\{a+b:a\in A,b\in B\}.$ We investigate diversities defined on $\mathbb{R}^{k}$ which are (Minkowski) linear [24]
(D5) $\delta(\lambda A)=\lambda\delta(A)\mbox{ and }\delta(A+B)=\delta(A)+\delta(B)$
and those which are (Minkowski) sublinear
(D6) $\delta(\lambda A)=\lambda\delta(A)\mbox{ and }\delta(A+B)\leq\delta(A)+\delta(B),$
for $\lambda\geq 0$ and $A,B$ nonempty finite subsets of $\mathbb{R}^{k}$. Many familiar diversities defined on $\mathbb{R}^{k}$ are Minkowski linear or sublinear (see below). We explore their properties and characterization.
@@
$h_{A}:\mathbb{R}^{k}\rightarrow\mathbb{R}:x\mapsto\sup\{a\cdot x:a\in A\}.$
Here $a\cdot x$ denotes the usual dot product in $\mathbb{R}^{k}$,
@@
We note that a set has the same support function as both its closure and convex hull.
We make use of the following properties of the support function, see [24, Chapter 1] for further details.
@@
    $h_{A+B}=h_{A}+h_{B}$ and $h_{\lambda A}=\lambda h_{A}$ for for non-empty, bounded $A,B$ and $\lambda\geq 0.$
@@
    If $A,B$ are nonempty convex compact sets then $A\subseteq B$ if and only $h_{A}(x)\leq h_{B}(x)$ for all $x\in\mathbb{R}^{k}$
@@
    A function $h:\mathbb{R}^{k}\rightarrow\mathbb{R}$ is the support function for some bounded nonempty set if and only if $h(x+y)\leq h(x)+h(y)$ and $h(\lambda x)=\lambda h(x)$ for all $x,y\in\mathbb{R}^{k}$ and $\lambda\geq 0$ (that is, $h$ is sublinear).
We often consider support functions restricted to $\mathbb{S}^{k-1}$, the unit sphere in $\mathbb{R}^{k}$, noting that a support function is determined everywhere by its values on $\mathbb{S}^{k-1}$. We note that the support function restricted to $\mathbb{S}^{k-1}$ of a nonempty set is bounded if and only if the set is bounded.
@@
    for a Borel measure $\nu$ on the sphere $\mathbb{S}^{k-1}$ satisfying
@@
    (Theorem 7) The extremal linear semidiversities are those where the support of $\nu$ is a finite, affinely independent set, which in turn correspond to a generalized circumradius (a Minkowski semidiversity) based on the simplex.
  3. 3.
    (Theorem 8) A diversity or semidiversity is sublinear if and only if it is the maximum of linear semidiversities (just like a function is convex if and only if it is the maximum of linear functions).
We then shift to studying the embeddings of finite diversities into linear or sublinear diversities. Questions regarding embeddings and approximate embeddings of metrics in normed spaces are central to metric geometry and its applications. Consider, for example, Menger’s characterizations of when a metric can be embedded in Euclidean space, or the vast literature applying metric embeddings to combinatorial optimizations (reviewed in [19, 8] and [16]).
@@
    (Theorem 11) A finite diversity can be embedded in a linear diversity if and only if it has negative type, meaning that
    $\sum_{A,B}x_{A}x_{B}\delta(A\cup B)\leq 0$
    for all vectors $x$ with zero sum and $x_{\emptyset}=0$.
@@
    (Theorem 12) A finite diversity can be embedded as a sublinear diversity if and only if it can be embedded in a Minkowski diversity (that is, a generalized circumradius) if and only if it is the maximum of a collection of negative type diversities.
@@ 2.1 Examples of Linear and Sublinear Diversities
We start with examples of diversities which are linear or sublinear. Note that for all diversities $(X,\delta)$ we have $\delta(\emptyset)=0$, even if that is not stated explicitly below.
@@
[…]
@@
    for finite $A\subseteq\mathbb{R}^{k}$. The diameter diversity is sublinear [24, pg 49].
@@
    $\delta_{1}(A)=\sum_{i=1}^{k}\max_{a,b\in A}(a_{i}-b_{i}).$
    for finite $A\subseteq\mathbb{R}^{k}$ [7]. For finite $A,B$ and $\lambda\geq 0$ we have
@@
    The circumradius of finite $A\subset\mathbb{R}^{k}$ with respect to the unit ball $\mathcal{B}$ is
    $\delta(A)=\min\{\lambda\geq 0:A\subseteq\lambda\mathcal{B}+x\mbox{ for some $x\in\mathbb{R}^{k}$}\}.$
    More generally, the Minkowski diversity $(\mathbb{R}^{k},\delta_{K})$ with kernel $K$ is equal to the generalized circumradius
    $\delta_{K}(A)=\inf\{\lambda\geq 0:A\subseteq\lambda K+x\mbox{ for some }x\in\mathbb{R}^{k}\},$
    for finite $A\subseteq\mathbb{R}^{k}$. Minkowski diversities are sublinear [3] but are not, in general, linear. For example, consider the circumradius diversity $(\mathbb{R}^{2},\delta)$. If $A=\{(0,0),(1,0)\}$ and $B=\{(0,0),(0,1)\}$ then $\delta(A+B)<\delta(A)+\delta(B)$.
    We assume that $K$ is closed, convex and has non-empty interior. We have elsewhere required that the kernel $K$ be bounded, however in this paper we will not require this. Note that if $K$ is an unbounded, $(\mathbb{R}^{k},\delta_{K})$ is a semidiversity rather than a diversity.
@@
    The mean-width diversity $(\mathbb{R}^{k},\delta_{w})$ is
[…]
@@
[…]
    That this is a sublinear diversity follows from the Minkowski inequality. See [13] Proposition 2.4, or [2] Proposition 10 in the case that $p=2$.
@@
    A zonotope $Z$ is a Minkowski sum of line segments and the length $\ell(Z)$ of the zonotope equals the sum of the length of the line segments. We define the zonotope diversity $(\mathbb{R}^{k},\delta_{z})$ where $\delta_{z}(A)$ is the minimum length of a zonotope containing $A$. We show that zonotope diversities are sublinear in Proposition 2.
In a Euclidean space $\mathbb{R}^{k}$, any non-negative linear combination of sublinear semidiversities is sublinear, and any non-negative linear combination of linear semidiversities is linear. Hence the set of sublinear semidiversities forms a convex cone, as does the set of linear semidiversities.

### v3

@@ Abstract
Diversities are an extension of the concept of a metric space which assign a non-negative value to every finite set of points, rather than just pairs. A general theory of diversities has been developed which exhibits many deep analogies to metric space theory but also veers off in new directions. Just as many of the most important aspects of metric space theory involve metrics defined on $\mathbb{R}^{k}$, many applications of diversity theory require a specialized theory for diversities defined on $\mathbb{R}^{k}$, as we develop here. We focus on two fundamental classes of diversities defined on $\mathbb{R}^{k}$: those that are Minkowski linear and those that are Minkowski sublinear. Many well-known functions in convex analysis belong to these classes, including diameter, circumradius and mean width. We derive surprising characterizations of these classes, and establish elegant connections between them. Motivated by classical results in metric geometry, and connections with combinatorial optimization, we then examine embeddability of finite diversities into $\mathbb{R}^{k}$. We prove that a finite diversity can be embedded into a linear diversity exactly when it is of negative type and that it can be embedded into a sublinear diversity exactly when it corresponds to a generalized circumradius.
@@ 1 Introduction
We start with a brief introduction to the theory of diversities, state our main results, and then give an overview of the paper.
A diversity BryantTupper12 is a pair $(X,\delta)$ where $X$ is a set and $\delta$ is a non-negative function defined on finite subsets of $X$ satisfying
@@
As such, diversities are set-based analogues of metric spaces. In fact, for a diversity $(X,\delta)$, if we let $d(x,y)=\delta(\{x,y\})$, then $(X,d)$ is a metric space, called the induced metric space BryantTupper12 .
Properties (D1) and (D2) are equivalent BryantNiesEtal21 to (D1) together with monotonicity
@@
Many well-known set functions are diversities: for any metric space the diameter of a set, the length of a connecting Steiner tree BryantTupper12 , and the length of a minimal traveling salesperson tour BryantTupper14 are diversities; within $\mathbb{R}^{k}$ the circumradius bryant2023diversities , the mean width BryantTupper14 , and the length of a smallest enclosing zonotope (see Proposition 2) are diversities. Two well-known set functions which fail to be diversities are average distance
$\pi(A)=\frac{1}{\binom{|A|}{2}}\sum_{\{a,b\}\subset A}d(a,b).$
which is not monotonic (D3), and volume of convex hull, which fails (D4).
There are broad classes of diversities just like there are broad classes of metrics. The $\ell_{1}$ metrics on $\mathbb{R}^{k}$ have the form
$d_{1}(a,b)=\sum_{i=1}^{k}|a_{i}-b_{i}|,$
while $\ell_{1}$ diversities on $\mathbb{R}^{k}$ BryantTupper14 have the form
$\delta_{1}(A)=\sum_{i=1}^{k}\max_{a,b\in A}|a_{i}-b_{i}|.$
For a finite set $X$, the class of (semi)-metrics of negative type is defined by the condition that $\sum_{a,b}x_{a}x_{b}\,d(a,b)\leq 0$ for all zero-sum vectors $x$. It includes $\ell_{1}$ metrics as a special case. In a similar way, for a finite set $X$, the class of semidiversities of negative type is defined by the condition that
$\sum_{A,B\subseteq X}x_{A}x_{B}\,\delta(A\cup B)\leq 0$ (1)
for all zero-sum vectors $x$ with $x_{\emptyset}=0$. It includes $\ell_{1}$ diversities as a special case. Metrics of negative type and diversities of negative type have multiple characterizations DezaLaurent97 ; WuBryantEtal19 . We note that WuBryantEtal19 only consider diversities of negative type, whereas we also consider semidiversities. We shall see (Proposition 8) that the conversion from one to the other is fairly automatic.
A Minkowski diversity $(\mathbb{R}^{k},\delta_{K})$ is determined by a compact convex set $K\subseteq\mathbb{R}^{k}$ with non-empty interior, known as the kernel bryant2023diversities . The diversity $\delta_{K}(A)$ of a finite subset $A\subseteq\mathbb{R}^{k}$ is the minimum amount that $K$ needs to be scaled so that a translate of $K$ covers $A$, that is, the generalized circumradius of $A$ with respect to $K$.
Our focus here is on the intersection of diversity theory, geometry and convex analysis. Recall that the Minkowski sum of two subsets $A,B\subseteq\mathbb{R}^{k}$ is given by $A+B=\{a+b:a\in A,b\in B\}$ and the scalar multiple of a set $A$ by $\lambda\in\mathbb{R}$ is given by $\lambda A=\{\lambda a:a\in A\}$. For $b\in\mathbb{R}^{k}$ we write $A+b$ as a shorthand for $A+\{b\}$. A semidiversity $(\mathbb{R}^{k},\delta)$ is Minkowski sublinear if it satisfies
(D5) $\delta(\lambda A)=\lambda\delta(A)\mbox{ and }\delta(A+B)\leq\delta(A)+\delta(B)$
for $\lambda\geq 0$ and all nonempty finite subsets $A,B\subseteq\mathbb{R}^{k}$. It is Minkowski linear if it also satisfies
(D6) $\delta(A+B)=\delta(A)+\delta(B)$
for all nonempty finite subsets $A,B\subseteq\mathbb{R}^{k}$. Many familiar diversities defined on $\mathbb{R}^{k}$ are Minkowski linear or sublinear (see below). We explore their properties and characterization. From here on we will use the terminology linear or sublinear as shorthand for Minkowski linear or sublinear respectively.
@@
$h_{A}:\mathbb{R}^{k}\to\mathbb{R}:x\mapsto\sup\{a\cdot x:a\in A\}.$
Here $a\cdot x$ denotes the usual scalar product in $\mathbb{R}^{k}$,
@@
We make use of the following properties of the support function, see (Schneider14, , Chapter 1) for further details.
@@
    A nonempty bounded set has the same support function as both its closure and convex hull.
@@
    $h_{A+B}=h_{A}+h_{B}$ and $h_{\lambda A}=\lambda h_{A}$ for non-empty, bounded $A,B$ and $\lambda\geq 0.$
@@
    If $A,B$ are nonempty convex compact sets then $A\subseteq B$ if and only if $h_{A}(x)\leq h_{B}(x)$ for all $x\in\mathbb{R}^{k}$.
We often consider support functions restricted to $\mathbb{S}^{k-1}$, the unit sphere in $\mathbb{R}^{k}$, noting that a support function is determined everywhere by its values on $\mathbb{S}^{k-1}$. We note that the support function restricted to $\mathbb{S}^{k-1}$ is bounded if and only if the set is bounded.
@@
    for a finite positive Borel measure $\nu$ on the sphere $\mathbb{S}^{k-1}$ satisfying
@@
    (Theorem 7) A diversity or semidiversity is sublinear if and only if it is the maximum of linear semidiversities (just like a function is convex if and only if it is the maximum of affine functions).
We then shift to studying the embeddings of finite diversities (that is, diversities $(X,\delta)$ with $X$ finite) into linear or sublinear diversities. We say that (semi)diversity $(X_{1},\delta_{1})$ can be embedded in (semi)diversity $(X_{2},\delta_{2})$ if there is a map $f\colon X_{1}\rightarrow X_{2}$ such that $\delta_{2}(f(A))=\delta_{1}(A)$ for all finite $A\subseteq X_{1}$. If the source $(X_{1},\delta_{1})$ is a diversity, then the embedding $f$ is automatically injective by $(D1)$, but in general we can embed semidiversities into diversities with non-injective $f$. Questions regarding embeddings and approximate embeddings of finite metrics in normed spaces are central to metric geometry and its applications. Consider, for example, Menger’s characterizations of when a metric can be embedded in Euclidean space, or the extensive literature applying metric embeddings to combinatorial optimization (reviewed in DezaLaurent97 ; matousek2013lectures ; indyk20178 ).
@@
    (Theorem 9) A finite semidiversity can be embedded in a linear diversity if and only if it can be embedded in a Minkowski diversity with simplex kernel (see Section 2.1 below) if and only if it is of negative type.
@@
    (Theorem 10) A finite semidiversity can be embedded in a sublinear diversity if and only if it can be embedded in a Minkowski diversity (see Section 2.1 below) if and only if it is the maximum of a finite collection of semidiversities of negative type on $X$.
That the characterization of linear embeddable diversities came out as tidy as Theorem 9 is surprising. Theorem 9 can, in many ways, be seen as the analogue of the celebrated and useful characterization of metrics of negative type as squares of Euclidean metrics. However the relationship between the metric and diversity characterizations is not as straightforward as it might first appear. The metric result is not a special case of the diversity result, in fact there are many metrics of negative type which are not the induced metrics of a diversity of negative type WuBryantEtal19 .
These results follow a general pattern which we have seen multiple times in the development of diversity theory. Sometimes the theory of diversities runs in parallel with that of metric spaces; other times it veers off in new directions. In the first paper on diversities BryantTupper12 the authors explored how concepts of hyperconvexity, injectivity and the tight span extended, and to an extent enriched, the analogous metric concepts. Diversities turn out to be an exemplary class of metric structures, exhibiting fascinating connections with model theory and Urysohn’s universal space BryantNiesEtal17 ; BryantNiesEtal21 ; Hallback20 ; hallback2020automorphism . Other directions that have been pursued are a diversity analogue of ultrametric and normed spaces HaghmaramNourouzi20 ; MehrabaniNourouzi20 ; haghmaram2022diversity and new diversity-based results in fixed point theory espinola2014diversities ; Piatek14 .
There are many reasons to believe that diversities are to hypergraphs what metrics are to graphs. The main contribution of BryantTupper14 is to show that the ‘geometry of graphs’ linking metric embeddings to approximation algorithms on graphs LinialLondonEtal95 has a parallel ‘geometry of hypergraphs’ linking diversity embeddings to approximation algorithms on hypergraphs. Jozefiak and Shepherd jozefiak2023diversity use this approach to obtain the best known approximation algorithms for several hypergraph optimization problems.
The approximation bounds achieved by LinialLondonEtal95 have been superseded by much tighter bound based on approximate embeddings of metrics of negative type. These tighter bounds motivate the study of negative type diversities WuBryantEtal19 . A geometric approach, as with metrics, has the potential to expand the metric embedding toolkit from graphs to hypergraphs. However diversity theory is new, and the fundamentals of the geometry of diversities are still being worked out. We see the results here as a contribution towards that endeavor.
The structure of the paper is as follows. In Section 2 we establish the basic properties and characterizations of linear and sublinear diversities. Subsection 2.1 gives examples of several linear and sublinear diversities, including defining Minkowski diversity. Subsection 2.2 establishes many basic properties of sublinear semidiversities that we need later for our main results. Subsection 2.3 gives Theorem 5, a complete characterization of linear semidiversities. Subsection 2.4 gives Theorem 7, our characterization of sublinear semidiversities. In Section 3 we switch to studying when finite diversities can be embedded in either linear or sublinear diversities. There we prove Theorem 9, a characterization of when a semidiversity is linear embeddable, and Theorem 10, a characterization of when a semidiversity is sublinear embeddable.
@@ 2.1 Examples of Linear and Sublinear Diversities
We start with examples of diversities which are linear or sublinear. Note that for all diversities $(X,\delta)$ we have $\delta(\emptyset)=0$, even if that is not stated explicitly below in the definitions.
@@
    Let $\|\cdot\|$ be any norm on $\mathbb{R}^{k}$. The diameter diversity (for this particular space) is given by
@@
    for finite $A\subseteq\mathbb{R}^{k}$. The diameter diversity on normed spaces is sublinear (Schneider14, , p. 49).
@@
    $\delta_{1}(A)=\sum_{i=1}^{k}\max_{a,b\in A}|a_{i}-b_{i}|$
    for finite $A\subseteq\mathbb{R}^{k}$ BryantTupper14 . For finite $A,B$ and $\lambda\geq 0$ we have (noting $\max_{a,b\in A}|a_{i}-b_{i}|=\max_{a,b\in A}(a_{i}-b_{i})$ )
@@
    The circumradius diversity of finite $A\subseteq\mathbb{R}^{k}$ with respect to the unit ball $\mathcal{B}$ is defined to be
    $\delta_{\mathcal{B}}(A)=\inf\{\lambda\geq 0:A\subseteq\lambda\mathcal{B}+x\mbox{ for some $x\in\mathbb{R}^{k}$}\}.$
    More generally, we define the Minkowski diversity $(\mathbb{R}^{k},\delta_{K})$ with kernel $K$ (also known as the generalized circumradius) to be
    $\delta_{K}(A)=\inf\{\lambda\geq 0:A\subseteq\lambda K+x\mbox{ for some }x\in\mathbb{R}^{k}\},$ (3)
    for finite $A\subseteq\mathbb{R}^{k}$. We assume that $K$ is compact and convex with non-empty interior. Minkowski diversities (including the circumradius) are sublinear bryant2023diversities but are not, in general, linear. For example, consider the circumradius diversity $(\mathbb{R}^{2},\delta_{\mathcal{B}})$. If $A=\{(0,0),(1,0)\}$ and $B=\{(0,0),(0,1)\}$ then $\delta_{\mathcal{B}}(A)=\delta_{\mathcal{B}}(B)=1/2$ but $\delta_{\mathcal{B}}(A+B)=\frac{1}{\sqrt{2}}$, so $\delta_{\mathcal{B}}(A+B)<\delta_{\mathcal{B}}(A)+\delta_{\mathcal{B}}(B)$.
@@
    The mean width diversity $(\mathbb{R}^{k},\delta_{w})$ BryantTupper14 is
    $\delta_{w}(A)=\frac{2}{\alpha_{k}}\int_{\mathbb{S}^{k-1}}h_{A}(x)\,\mathrm{d}\nu(x)$
    where $\nu(x)$ is the (uniform) Haar measure on the sphere and $\alpha_{k}$ is chosen so that $\delta_{w}(\{a,b\})=\|a-b\|_{2}$ for all $a,b\in\mathbb{R}^{k}$. The value $\delta_{w}(A)$ is proportional to the mean width of the convex hull of $A$. Mean width diversities are linear (Schneider14, , p. 50) and their induced metric is the Euclidean metric.
@@
    $\delta^{(p)}_{w}(A)=\frac{1}{\alpha_{k}}\left[\int_{\mathbb{S}^{k-1}}|w_{A}(x)|^{p}\,\mathrm{d}\nu(x)\right]^{1/p}.$
    That this is a sublinear diversity follows from the Minkowski inequality. See (HaghmaramNourouzi20, , Prop. 2.4), or (BryantCioica-LichtEtal21, , Prop. 10) in the case that $p=2$.
@@
    A zonotope $Z$ is a Minkowski sum of line segments. By linearity, the mean width diversity of a zonotope equals the sum of lengths of the line segments it is formed from (see joos2023isoperimetric ). This sum is called the length $\ell(Z)$ of the zonotope, and is well-defined even though a zonotope can be expressed as the sum of multiple different collections of line segments.
    We define the zonotope diversity $(\mathbb{R}^{k},\delta_{z})$, where $\delta_{z}(A)$ is the minimum length of a zonotope containing $A$, and show that it is a sublinear diversity in Proposition 2. Algorithms for computing the length of the minimum enclosing zonotope can be found in guibas2003zonotopes .